# Hybrid Near-field and Far-field Localization with Holographic MIMO

### v1

@@ III-B2 Scatter Location Estimation
$s.t.$ |$|\{i|\bm{v}_{i}\neq\bm{0}\}|\leq L_{max},$ (23b)
where $\bm{v}_{i}=[u_{i1}^{\bm{Q}},...,u_{iK}^{\bm{Q}}]\in\mathbb{C}^{1\times K}$. Since the numbers of scatters are generally much smaller than the number of atom channels[23], we can use the sparse recovery method[20] to estimate the sparse vector $\bm{u}^{\bm{Q}}_{k}$[22]. We define the support set $\mathcal{X}$ as the indices of the atom signals corresponding to the estimated scatter locations. In each iteration, a new index is added to the support set. Note that the atom signal for each user is different, the residual signal for different users is approximated using their corresponding atom signals, and then the location with the minimal modulus sum of the new residual signals is selected as the estimated location of a scatter. Mathematically, we have
$i=\arg\max_{i}\sum_{k=1}^{K}|(\bm{\Lambda}^{k}_{i})^{H}\bm{e}^{(c)}_{k}|$ (24)
@@ IV RIS Phase Shift Optimization
In this section, we first formulate the RIS phase shift optimization problem and propose an optimization algorithm to solve it.
@@ IV-A Phase Shift Optimization Problem Formulation
The selected optimization metric is CRB, a standard metric for assessing the parameter estimation error. The CRB is derived from the model of the received signal for $c+1$ cycles and given by
@@
Specifically, we formulate the optimization problem to minimize the sum of weighted CRBs of the range and angle estimation errors. Note that we estimate the angles and ranges for NF users, while only the angles for FF users. In other words, the location parameters for NF and FF users are different. To this end, we give the expressions of the weighted CRBs for NF and FF users separately.
@@ IV-A1 CRBs for NF Users
If the estimated location $\bm{p}_{k}^{(c)}$ is in the NF region in the $c$-th cycle, we expect that in the $(c+1)$-th cycle, the range and angles of the $k$-th user are estimated, and the estimated parameters are denoted by $\bm{p}_{k}^{(c+1)}=[R_{k},\theta_{k},\phi_{k}]^{T}$. The expressions of CRBs for the NF case are provided in proposition IV-A1.
@@ IV-A3 Optimization Problem Formulation
where $\bm{J}_{NF,k}$ or $\bm{J}_{FF,k}$ is the Fisher information matrix of $\bm{p}_{k}$ when the $k$-th user is located in the NF or FF region and is a function of the RIS phase shifts $\bm{\beta}^{(c+1)}$. Here $\bm{W}_{NF}=\text{diag}\{w_{1},w_{2},w_{3}\}$ and $\bm{W}_{FF}=\text{diag}\{w_{4},w_{5}\}$ are the weight matrices of the CRBs in the NF and FF regions, respectively. $\Psi_{NF}$ and $\Psi_{FF}$ are the sets of estimated NF and FF users, which are mathematically given by
@@ IV-B RIS Phase Shift Optimization Algorithm
We design an optimization algorithm based on the CCM method to tackle (P2). Due to the constant modulus constraint (41b), the optimization problem (P2) is non-convex, which is numerically difficult to handle. Fortunately, the solution can be considered as lying on the CCM to satisfy the constant modulus, where the manifold is represented as
@@
The main idea of the CCM-based optimization method is to iteratively apply gradient descent in the manifold space. After several iterations, the algorithm terminates when the difference between two iterations of $f(\bm{P},\bm{\beta}^{(c+1)})$ is less than a constant $\zeta$. The CCM-based optimization method is composed of four main steps in each iteration:
@@ IV-B4 Retract onto the manifold
2:  while $|g(\bm{\beta}^{(c+1)}_{j+1})-g(\bm{\beta}^{(c+1)}_{j})|<\zeta$ do
@@ V-B Localization Performance
Localization errors primarily arise from two sources: one is the failure to estimate the grid closest to the user, which we refer to as grid misjudgment; the other is the error between the user’s continuous location in actual space and the discrete grids, known as grid mismatch, reflecting the inherent limitation of discretizing continuous space[29].
Mathematically, suppose the real location of the $k$-th user is $\bm{p}_{k}$, the estimation result is $\bm{z}_{m^{\prime}}$, and the nearest grid for the $k$-th user is $\bm{z}_{m}$. Then the average localization error can be given by
@@
where $l_{e}=\|\bm{p}_{k}-\bm{z}_{m^{\prime}}\|$ is the localization error. The first term corresponds to grid mismatch, and the second term is grid misjudgment. In the following, we analyze these two sources of errors respectively.
@@ V-B1 Number of Cycles
where $\mathcal{R}_{k,m}$ is the decision region for the $m$-th candidate location. The integral in (50) is the probability that the estimated location is the $m$-th candidate location, given the ground truth that the location of the $k$-th user is $\bm{p}_{k}$. Here $\gamma_{m}^{k}$ is the error parameter, defined as $\gamma_{m}^{k}=\|\bm{p}_{k}-\bm{z}_{m}\|$.
@@
$\mathcal{R}_{k,m}=\{\bm{g}_{k}^{(C)}:|\bm{g}_{k}^{(C)}-\bm{\mu}_{m}|^{2}\leq|\bm{g}_{k}^{(C)}-\bm{\mu}_{m^{\prime}}|^{2},\forall m^{\prime}\neq m\},$ (52)
@@
$\mathcal{R}_{k,m}=$$\{\bm{g}_{k}^{(C)}:|\bm{\xi}_{k}|^{2}\leq|\bm{\xi}_{k}+\bm{\mu}_{m}-\bm{\mu}_{m^{\prime}}|^{2},\forall m^{\prime}\neq m\},$
@@
$\lim_{C\to\infty}\mathbb{P}(\bm{g}_{k}^{(C)}\in\mathcal{R}_{k,m})=1.$ (54)
@@
When the user is not at a grid, if we assume the midpoint of $\bm{z}_{m}$ and $\bm{z}_{m^{\prime}}$ also corresponds the midpoint of $\bm{\mu}_{m}$ and $\bm{\mu}_{m^{\prime}}$, then we have $\mathbb{E}(\bm{\xi}_{k})\leq|\bm{\mu}_{m}-\bm{\mu}_{m^{\prime}}|/2$. Hence, the $|\bm{\mu}_{m}-\bm{\mu}_{m^{\prime}}|^{2}+2\bm{\xi}_{k}(\overline{\bm{\mu}_{m}-\bm{\mu}_{m^{\prime}}})\geq 0$ is still alwasy true, and we still have $\lim_{C\to\infty}\mathbb{P}(\bm{g}_{k}^{(C)}\in\mathcal{R}_{k,m})=1$. However, the number of cycles does not affect the grid mismatch error. Therefore, the localization error would converge to a non-zero number.
@@ VI Simulation Results
In this section, we present simulation results demonstrating the performance of the proposed method. Simulations are conducted on a personal computer with a 1.80GHz Intel Core i7 processor. The RIS is placed on the plane $x=0$, and the center of RIS is at the origin $(0,0,0)$m. We use an RIS with $10\times 10$ elements, and the element spacing is $\lambda/2=0.005$m. The center transmit frequency is 28GHz, and the frequency spacing between different users is set as $\Delta f=120$kHz[33]. The localization range is $10$m in radius. The azimuth and elevation angles are both uniformly sampled with sampling spacing $\pi/10$, and the range is sampled with spacing $0.25$m. The number of cycles is set as $20$. We conduct $T=1000$ independent trials to obtain the average result. The simulation parameters are listed in Table I unless otherwise stated.
TABLE I: Simulation Parameters
@@
Transmit center frequency $f_{c}$ $28$GHz
Frequency spacing $\Delta f$ $120$KHz
@@
VI-A Performance Comparison
Fig. 4 shows the RMSE of angle or range estimation versus the signal-to-noise ratio (SNR) when the user is in the FF or the NF region. The SNR is defined as $P_{s}/\sigma^{2}$, where $P_{s}=((\bm{h}^{A}_{k})^{T}\bm{I}\bm{h}_{k0}^{t})^{2}$ is the received signal power of the direct path when the phase shifts of all the RIS elements are $1$. The proposed method is compared with two kinds of comparison methods.
  • First, the results of a RIS-enabled NF localization method (AML in [16]) and a RIS-enabled FF localization method (ANM in [34]) are provided for comparison to reveal the advantage of the proposed algorithm.
  • Second, the proposed algorithm is combined with pure NF or FF models to show the gain brought by the hybrid model, labeled as Proposed (NF) and Proposed (FF).
@@ VI-A Performance Comparison
Fig. 4 (a) shows the angle estimation RMSE of the proposed algorithm and other comparison algorithms when the user is located in the FF region. Compared with AML and ANM, it shows that the proposed method outperforms all other comparison methods for FF user localization. We also observe that for algorithms other than ANM, as SNR increases, the RMSE of the angle estimation first decreases and then remains fixed. The RMSE does not decrease when the SNR is very large because the estimated location must be at one of the sampled grids, while the actual location of the users could be off-grid. The ANM does not suffer from the grid mismatch problem since it is an off-grid algorithm.
@@
Compared with the results of pure NF or FF models, Fig. 4 (a) reveals that the RMSEs of the proposed algorithm when applied to the hybrid model and FF model are similar, indicating the hybrid model behaves similarly for FF user localization although with the interference of NF candidate locations. We also observe that the proposed algorithm performs worse in the NF model than in the hybrid model for FF user localization. This is because, for the pure NF model, the range of the FF region is also sampled, resulting in more candidate locations in the FF region and making the algorithm more difficult to converge. At the same time, the channels of the sampled FF candidate locations in the NF model have strong correlations, which leads to energy dispersion especially when estimating the scatters locations, resulting in lower estimation performance. We conducted more simulations in the following to support the above arguments.
Fig. 4 (b) shows the angle estimation RMSE of the proposed algorithm when the user is located in the NF region. It shows that the proposed method provides higher localization accuracy than the compared algorithms. We also observe that the FF model has a higher RMSE than the hybrid model when estimating the NF angles. This is due to the model mismatch of the FF model when describing the NF channels. Therefore, the hybrid model has the advantage of higher localization accuracy compared to the pure FF model. Furthermore, for NF user localization, the angle estimation result of the proposed algorithm applied to the hybrid and NF models is similar.
@@
Fig. 4 (c) shows the range estimation accuracy for NF users versus SNR. For range estimation, the proposed method outperforms AML by more than 55% for all SNR conditions. It shows that the range estimation result of the NF model is not as good as those of the hybrid model for NF user localization, which reveals the superiority of the hybrid model.
Figure 5: Angle estimation performance versus number of cycles for FF users.
Fig. 5 illustrates the angle estimation RMSE versus the number of cycles for FF users. It shows that for the hybrid model and FF model, the estimation performance converges after $C=30$ cycles, while for the NF model, the result remains fixed after $C=35$ cycles, which proves the argument that the NF model is harder to converge.
Figure 6: Angle estimation performance versus localization ranges for FF users.
Fig. 6 shows the angle estimation RMSE versus the localization range. We observe that as the localization range expands, the RMSE under the NF model increases, while the RMSE remains unchanged for the hybrid model. This is because as the localization range expands, for the NF model, the number of candidate locations also increases, while they remain the same for the FF and hybrid model. Also, note that for a sufficiently small estimation range, the NF model could outperform the hybrid model. This is due to that the NF model is more accurate and the FF model is the approximation of the NF model for the FF region. When the localization range is sufficiently small, the above issue of the NF model is mitigated, and the advantage of the NF model being more accurate exceeds the losses caused by the above problem.
Figure 7: CPU time of the proposed algorithm versus number of cycles.
Fig. 7 reveals the CPU time versus number of cycles for three different models. The results indicate that the proposed algorithm has the lowest complexity when applied to the FF model, followed by the hybrid model, and the highest complexity in the NF model. This is due to that the FF model has the fewest candidate locations, while the NF model has the most. This shows the advantage of low algorithm complexity of the hybrid model over the NF model.
In conclusion, the proposed method outperforms other RIS-based localization algorithms in both angle and range estimation accuracy. Compared with different models, the hybrid model offers superior localization accuracy compared to the pure FF model and has the dual advantage of higher localization accuracy and lower algorithm complexity compared to the pure NF model.
@@ VI-B Localization Accuracy and Complexity Trade-off
Figure 8: Average angle estimation performance of the NF and FF users for different angle sampling spacing.
Figure 9: CPU time of the proposed algorithm versus angle sampling spacing.
This subsection examines the effects of hyperparameters on localization accuracy and algorithm complexity. The analysis in Sec. V shows that the sampling spacing, the number of cycles, and the number of RIS elements are the most important factors for both.

### v2

@@ III-B2 Scatter Location Estimation
$s.t.$ |$\left|\{i|\bm{v}_{i}\neq\bm{0}\}\right|\leq L_{max},$ (23b)
where $\bm{v}_{i}=[u_{i1}^{\bm{Q}},...,u_{iK}^{\bm{Q}}]\in\mathbb{C}^{1\times K}$. Since the numbers of scatters are generally much smaller than the number of atom channels[30], we can use the sparse recovery method[27] to estimate the sparse vector $\bm{u}^{\bm{Q}}_{k}$[29]. We define the support set $\mathcal{X}$ as the indices of the atom signals corresponding to the estimated scatter locations. In each iteration, a new index is added to the support set. Note that the atom signal for each user is different, the residual signal for different users is approximated using their corresponding atom signals, and then the location with the maximum correlation with the residual signal is selected as the estimated location of a scatter. Mathematically, we have
$i=\arg\max_{i}\sum_{k=1}^{K}\left|(\bm{\Lambda}^{k}_{i})^{H}\bm{e}^{(c)}_{k}\right|$ (24)
@@ IV RIS Phase Shift Optimization
In this section, we first formulate the RIS phase shift optimization problem and propose an algorithm to solve it.
@@ IV-A Phase Shift Optimization Problem Formulation
The selected optimization metric is CRB, a standard metric for assessing the parameter estimation error[31]. The CRB is derived from the model of the received signal for $c+1$ cycles, which is given by
@@
Specifically, we formulate the optimization problem to minimize the sum of weighted CRBs of range and angle estimation errors. We introduce a weight matrix to address the different unit of angle and range CRBs [32]. Note that we estimate angles and ranges for NF users, while only angles for FF users. Hence, we give the expressions of the weighted CRBs for NF and FF users separately.
@@ IV-A1 CRBs for NF Users
If the estimated location $\bm{p}_{k}^{(c)}$ is in the NF region in the $c$-th cycle, we expect that in the $(c+1)$-th cycle, the range and angles of the $k$-th user are estimated, denoted by $\bm{p}_{k}^{(c+1)}=[R_{k},\theta_{k},\phi_{k}]^{T}$. The expressions of CRBs for the NF case are provided in proposition IV-A1.
@@ IV-A3 Optimization Problem Formulation
where $\bm{J}_{NF,k}$ or $\bm{J}_{FF,k}$ is the Fisher information matrix of $\bm{p}_{k}$ when the $k$-th user is located in the NF or FF region and is a function of the RIS phase shifts $\bm{\beta}^{(c+1)}$. Here $\bm{W}_{NF}=\text{diag}\{w_{1},w_{2},w_{3}\}$ and $\bm{W}_{FF}=\text{diag}\{w_{4},w_{5}\}$ are the weight matrices of the CRBs in the NF and FF regions, respectively [32]. $\Psi_{NF}$ and $\Psi_{FF}$ are the sets of estimated NF and FF users, which are mathematically given by
@@ IV-B RIS Phase Shift Optimization Algorithm
We design an optimization algorithm based on the complex circle manifold (CCM) method to tackle (P2). Due to the constant modulus constraint (41b), the problem (P2) is non-convex, which is numerically difficult to handle. Fortunately, the solution can be considered as lying on the CCM to satisfy the constant modulus, where the manifold is represented as
@@
The main idea of the CCM-based optimization method is to iteratively apply gradient descent in the manifold space. After several iterations, the algorithm terminates when the difference between two iterations of $f(\bm{P},\bm{\beta}^{(c+1)})$ is less than a constant $\zeta$ or the number of iterations exceeds $I$, where $\zeta$ and $I$ are selected to comply with the time constraints of the optimization step. The CCM-based optimization method is composed of four main steps in each iteration:
@@ IV-B4 Retract onto the manifold
2:  while $|g(\bm{\beta}^{(c+1)}_{j+1})-g(\bm{\beta}^{(c+1)}_{j})|>\zeta$ and $j<I$ do
@@ V-B Localization Performance
Localization errors primarily arise from two sources: one is the failure to estimate the grid closest to the user, which we refer to as grid misjudgment; the other is the error between the user’s continuous location in actual space and the discrete grids, known as grid mismatch, reflecting the inherent limitation of discretizing continuous space[37]. Mathematically, suppose the real location of the $k$-th user is $\bm{p}_{k}$, the estimation result is $\bm{z}_{m^{\prime}}$, and the nearest grid for the $k$-th user is $\bm{z}_{m}$. Then the average localization error can be given by
@@
where the first term corresponds to grid mismatch, and the second term is grid misjudgment. In the following, we analyze these two sources of errors respectively.
@@ V-B1 Number of Cycles
where $\mathcal{R}_{km}$ is the decision region for the $m$-th candidate location. The integral in (50) is the probability that the estimated location is the $m$-th candidate location, given the ground truth that the location of the $k$-th user is $\bm{p}_{k}$. Here $\gamma_{m}^{k}$ is the error parameter, defined as $\gamma_{m}^{k}=\|\bm{p}_{k}-\bm{z}_{m}\|$.
@@
$\mathcal{R}_{km}\!=\!\{\bm{g}_{k}^{(C)}\!:\!|\bm{g}_{k}^{(C)}\!-\!\bm{\mu}_{m}|^{2}\!\leq\!|\bm{g}_{k}^{(C)}\!-\!\bm{\mu}_{m^{\prime}}|^{2}\!,\forall m^{\prime}\!\neq\!m\},$ (52)
@@
$\mathcal{R}_{km}=$$\{\bm{g}_{k}^{(C)}:|\bm{\xi}_{k}|^{2}\leq|\bm{\xi}_{k}+\bm{\mu}_{m}-\bm{\mu}_{m^{\prime}}|^{2},\forall m^{\prime}\neq m\},$
@@
$\lim_{C\to\infty}\mathbb{P}(\bm{g}_{k}^{(C)}\in\mathcal{R}_{km})=1.$ (54)
@@
When the user is not at a grid, if we assume the midpoint of $\bm{z}_{m}$ and $\bm{z}_{m^{\prime}}$ also corresponds the midpoint of $\bm{\mu}_{m}$ and $\bm{\mu}_{m^{\prime}}$, then we have $\mathbb{E}(\bm{\xi}_{k})\leq|\bm{\mu}_{m}-\bm{\mu}_{m^{\prime}}|/2$. Hence, the $|\bm{\mu}_{m}-\bm{\mu}_{m^{\prime}}|^{2}+2\bm{\xi}_{k}(\overline{\bm{\mu}_{m}-\bm{\mu}_{m^{\prime}}})\geq 0$ is still alwasy true, and we still have $\lim_{C\to\infty}\mathbb{P}(\bm{g}_{k}^{(C)}\in\mathcal{R}_{km})=1$. However, the number of cycles does not affect the grid mismatch error. Therefore, the localization error would converge to a non-zero number.
@@ VI Simulation Results
In this section, we present simulation results demonstrating the performance of the proposed method. The RIS is placed on the plane $x=0$, and its center is at the origin $(0,0,0)$m. We use an RIS with $10\times 10$ elements and the element spacing is $\lambda/2=0.03$m. The localization range is $10$m in radius [41]. The NF range for the considered RIS is approximately $4.9$m, while the rest region is FF. The users are equally distributed in NF and FF regions. We consider a multipath environment with $3$ randomly distributed scatters. The azimuth and elevation angles are both uniformly sampled with spacing $\pi/10$, and the range is sampled with spacing $0.25$m. The center transmit frequency is 5GHz [42]. The number of cycles is set as $20$. We conduct $T=1000$ independent trials to obtain the average result. The simulation parameters are listed in Table II unless otherwise stated. The proposed method is compared with following methods.
  • Approximate maximum likelihood (AML) scheme[22]: a RIS-enabled NF localization method based on maximum likelihood. This method decouples the angle and range parameters and iteratively searches the angle and range to fit best with the received signal.
  • Atomic norm minimization (ANM) scheme[43]: an off-grid direction of arrival estimation method based on atomic norm that targets FF users localization.
  • NF scheme: The proposed algorithm is combined with pure NF model to show the gain brought by the hybrid model, labeled as NF.
  • FF scheme: The proposed algorithm is combined with pure FF model, labeled as FF.
For all compared algorithms, we employ the same system setup and localization protocol. For NF methods, including AML and the proposed method with the NF model, the whole localization range is treated as in the NF, i.e., the range of FF is also sampled, while for FF methods, we only sample the angle.
TABLE II: Simulation Parameters
@@
Transmit center frequency $f_{c}$ $5$GHz
@@
VI-A Performance Evaluation
@@
Fig. 4 shows the RMSE of angle or range estimation versus the signal-to-noise ratio (SNR) for NF or FF users. The SNR is defined as $P_{s}/\sigma^{2}$, where $P_{s}=((\bm{h}^{A}_{k})^{T}\bm{I}\bm{h}_{k0}^{t})^{2}$ is the received signal power of the direct path when the phase shifts of all the RIS elements are $1$. Fig. 4 (a) and (b) show the angle estimation RMSE of the proposed algorithm and other comparison algorithms for FF and NF users, respectively. They indicate that the proposed method outperforms all other methods for FF and NF user localization. We also observe that for algorithms other than ANM, as SNR increases, the RMSE of the angle estimation first decreases and then remains fixed. The RMSE does not decrease when the SNR is very high because the estimated location must be at one of the sampled grids, while the actual location of the users is off-grid. The ANM does not suffer from the grid mismatch problem since it is an off-grid algorithm.
@@
Fig. 4 (a) also reveals the RMSEs of the proposed algorithm with the hybrid model and FF model are similar, indicating that the hybrid model can achieve the same accuracy for FF user localization although with the interference of NF candidate locations. We also observe that the proposed algorithm performs worse in the NF model than in the hybrid model for FF users. This is because, for the NF model, the range of the FF region is also sampled, resulting in more candidate locations in the FF region and making the algorithm more difficult to converge. At the same time, the channels of the sampled FF candidate locations in the NF model have strong correlations, which leads to energy dispersion especially when estimating the scatters locations. In Fig. 4 (b), we observe that the FF model has a higher RMSE than the hybrid model when estimating the NF angles. This is due to the model mismatch of the FF model when describing the NF channels.
@@
Fig. 4 (c) shows the range estimation accuracy for NF users versus SNR. For range estimation, the proposed method outperforms AML by more than 55% for all SNR conditions. It shows that the range estimation result of the NF model is not as good as those of the hybrid model for NF user localization, which reveals the superiority of the hybrid model. Additionally, the improvement of the hybrid model over the NF model is bigger in range compared to angle estimation. The reason for this phenomenon is that the NF model samples the range and angle in the whole region, i.e., the range of FF is also sampled. However, the steering vectors of the different ranges in the FF are very similar. If a NF user is misestimated as a FF point, the resulting range error becomes significantly larger, leading to an overall degradation in estimation accuracy.
In Fig. 4, we can observe that the proposed method significantly outperforms both AML and ANM. The performance gain can be attributed to the following reasons: Firstly, neither AML nor ANM effectively handles the scattering paths, and the interference of scattering paths reduces localization accuracy. Secondly, AML decouples angle and range estimation and updates them iteratively. While this approximated method reduces algorithm complexity, it also impacts performance. Moreover, ANM is a FF localization method and employs plane wave model, which inaccurately describes NF signals.
Figure 5: Angle estimation performance for different localization scenarios.
In the following, the angle sampling spacing is set as $\pi/20$ to reduce grid mismatch. Fig. 5 shows the influence of the number of users and scatters. We can observe that increasing the number of users degrades angle estimation accuracy, as the RIS beamforming becomes less precise when simultaneously serving multiple users. Moreover, increasing the number of scatters reduces localization accuracy due to interference from multipath propagation. Furthermore, by comparing the results obtained directly using the OMP algorithm, i.e., the first step of the localization algorithm (labeled as coarse) with those obtained through the three-step localization algorithm (labeled as refined), it is evident that the scattering path elimination technique can improve the localization accuracy.
Fig. 6 shows the angle estimation performance versus the scattering path power. The CRB is calculated given true user and scatter locations. It is observed that for low scattering path power and dense angle sampling, the estimation RMSE can approach the CRB. However, the effectiveness of scattering path elimination becomes limited for high scattering path interference, causing the RMSE to deviate from the CRB. Additionally, sparse angle sampling exacerbates the grid mismatch problem, further leading to accuracy degradation.
Figure 6: Angle estimation performance for different interference power.
@@ VI-B Localization Accuracy and Complexity Trade-off
This subsection examines the effects of hyperparameters on localization accuracy and algorithm complexity. We focus on the sampling spacing, the number of cycles, and the number of RIS elements, as they are the most important factors for both according to the analysis in Sec. V.